OpenStudy (verarae):

11. The coordinates of endpoint A are A(5, -15). The midpoint of AC is B(5, -8). Find the coordinate of endpoint C. (1 point work, 1 point x2, one point y2) 12. Find the distance between A(12, 6) and B(-8, 18). Leave your answer in simplest radical form. (1 point work, 1 point answer) pretty please help me with this . 13. Find the area of a circle with diameter 8 in. Round to the nearest tenth. (1 point answer, 1 point units)

OpenStudy (optyprompty):

────────▌▒█───────────▄▀▒▌ ────────▌▒▒▀▄───────▄▀▒▒▒▐ ───────▐▄▀▒▒▀▀▀▀▄▄▄▀▒▒▒▒▒▐ ─────▄▄▀▒▒▒▒▒▒▒▒▒▒▒█▒▒▄█▒▐ ───▄▀▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▀██▀▒▌ ──▐▒▒▒▄▄▄▒▒▒▒▒▒▒▒▒▒▒▒▒▀▄▒▒▌ ──▌▒▒▐▄█▀▒▒▒▒▄▀█▄▒▒▒▒▒▒▒█▒▐ ─▐▒▒▒▒▒▒▒▒▒▒▒▌██▀▒▒▒▒▒▒▒▒▀▄▌ its raining Doges... ─▌▒▀▄██▄▒▒▒▒▒▒▒▒▒▒▒░░░░▒▒▒▒▌ ─▌▀▐▄█▄█▌▄▒▀▒▒▒▒▒▒░░░░░░▒▒▒▐ ▐▒▀▐▀▐▀▒▒▄▄▒▄▒▒▒▒▒░░░░░░▒ ▐▒▒▒▀▀▄▄▒▒▒▄▒▒▒▒▒▒░░░░░░▒▒▒▐ ─▌▒▒▒▒▒▒▀▀▀▒▒▒▒▒▒▒▒░░░░▒▒▒▒▌ ─▐▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▐ ──▀▄▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▄▒▒▒▒▌ ────▀▄▒▒▒▒▒▒▒▒▒▒▄▄▄▀▒▒▒▒▄▀ ───▐▀▒▀▄▄▄▄▄▄▀▀▀▒▒▒▒▒▄▄▀ ──▐▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▀

OpenStudy (mhchen):

For 11. Find the distance between A and B using the distance formula. I assume you learned that? Then since B is the midpoint, the distance from A to B is the same as the distance from B to C. 12. Use the distance formula. (Or the pythagorean theorem). 13. Area of a circle is radius^2 * pi

OpenStudy (ben00):

Boy there's a lot of people!

OpenStudy (ben00):

Also can I get a medal for nothing? I mean the doge meme got one.

OpenStudy (optyprompty):

Lol.

OpenStudy (optyprompty):

I didn't know it would get a medal.

OpenStudy (ben00):

Someone just medal me. :/

OpenStudy (1davey29):

11) The distance from the midpoint (B) and the endpoint (A) is 7 units, as the x is the same for both, so you only deal with the y values. The distance from A to B is also the same as the distance from B to C, as it's a midpoint, so it would be 7 units in the direction from A to B, or right, which would just be -8+7 which is -1, so C would be (5,-1) 12) Here, the distance formula should be applied (sqrt((x2-x1)^2+(y2-y1)^2)). Plugging in your values, you would get sqrt((12-(-8))^2+(6-18)^2), which simplifies to sqrt(544) or 4*sqrt(34) 13) Area of a circle is pi*r^2. r=d/2, which is 4 in. in this case. A=pi*4^2=pi*16=50.3 in.